The product of two consecutive integers is 4 less than 4 times their sum. Find the integers.
Answer by edjones(8007)

Let the integers be n and n+1
n(n+1)=4(n+n+1)-4
n^2+n=4(2n+1)-4
n^2+n=8n+4-4
n^2-7n=0
n(n-7)=0
n=0, n=8
.
There are 2 correct answers:
0 and 1
8 and 9
.
Ed